# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v2

@@ 6 Discussion
It is important to consider the assumptions and limitations of the proposed approach. One of the main drawbacks of the model is the additivity assumption imposed on the physical, domain and class components of the response. It is expected that the model will perform sub-optimally when this assumption is violated. Furthermore, the accuracy of the inferred physics-grounded latent variables will depend on the relative contribution of the physics, domain and class influences to the measured response. Significant domain and class contributions to the response, or violating the additivity assumption of Equation 2.3, can lead to inaccurate inference of physics-grounded latent variables and large uncertainty in the predictions. Additionally, the model requires that multiple types of data are available, namely measurements of the structural response and information on domain and class. In SHM applications, this might necessitate data alignment procedures of response measurements, environmental conditions and damage level descriptions, and could potentially limit the immediate applicability of the proposed architecture. It is not possible to provide a rule-of-thumb about the amount of data required for effective training. This would be dependent on the specific problem, noise levels, physics-based model, and accuracy of the domain and class information, and also on the particular architecture choices (e.g. the number and depth of layers used in the feed-forward NNs in the encoder and decoder). In some applications, the incorporation of the known physics might lead to a reduction in the data requirements. This has not been investigated in the current paper, since it is believed that it would be strongly dependent on the case study chosen, rather than offering any general insight. It is worth mentioning that the interaction between the encoder, decoders, the GRL, and the known physics can be unintuitive in some applications, limiting the applicability of the approach and potentially necessitating implicit supervision by a human expert. Finally, it is known that adversarial training can be unstable [Wiatrak2020]. Throughout this work, occasional instability and overfitting were observed when using small datasets and large batch sizes. Depending on the case study and the value of the $\lambda$ hyperparameter, oscillatory behavior may also occur. We found that these issues could be addressed by adjusting the $\lambda$ hyperparameter, implementing early stopping based on the value of the ELBO on a held-out validation set, and reducing the batch size.
@@ 7 Conclusions
The present work contributes to the emerging applications of probabilistic generative models in engineering, by investigating disentangled and invariant representation learning as a tool for grounding VAE to the known physics. Specifically, a Physics-Enhanced Machine Learning strategy utilizing a VAE architecture is proposed, with the aim of learning a disentangled representation of physical, domain and class confounding influences that are present in the response measurements of physical systems. This is achieved by having the encoder, decoder and latent space of the VAE be semantically and functionally separated into data-driven and physics-grounded branches. An easy to implement regularization method based on the GRL is used to constrain the data-driven components, resulting in a model that preferentially utilizes the known physics. An interpretable and intuitive hyperparameter is used to specify the strength of GRL, and wether the model is trained in a collaborative or adversarial manner. Moreover, a strategy for quantifying the type and relative amount of information encoded in different sets of latent variables is proposed, yielding insights on the degree of disentanglement achieved by the model.
@@
The code and data (generated via the synthetic use cases) needed to replicate the results shown in this paper can be accessed through the link: https://doi.org/10.5281/zenodo.15813028.
@@ Appendix A Implementation details
The encoder reduces the dimensionality of the input measurements and maps them to vectors of mean values $\bm{\mu}_{\mathrm{\phi}}(\bm{x})$, standard deviations $\bm{\sigma}_{\mathrm{\phi}}(\bm{x})$ and a lower triangular matrix $\bm{L}^{\prime}_{\mathrm{\phi}}(\bm{x})$. The lower triangular factor $\bm{L}_{\mathrm{\phi}}(\bm{x})=\bm{L}^{\prime}_{\mathrm{\phi}}(x)+\bm{\sigma}_{\mathrm{\phi}}(\bm{x})\bm{I}$ is the Cholesky decomposition factor of the covariance matrix $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})$, i.e. $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})=\bm{L_{\mathrm{\phi}}(\bm{x})L_{\mathrm{\phi}}(\bm{x})}^{T}$, such that the posterior distribution corresponding to each input is a multivariate Normal distribution. The encoder outputs the log of the standard deviations, which are then exponentiated to avoid negative values. The reparametrization trick [Kingma2022] is exploited to define the latent variables $\bm{z}$ as a deterministic transformation of a noise variable $\bm{\epsilon}\sim p(\bm{\epsilon})$. This facilitates the computation of unbiased Monte Carlo gradient estimates of the objective with respect to the variational parameters, using automatic differentiation. With the exception of the introductory examples presented in Section 2.4, the encoder is everywhere formulated as a single feed-forward NN using a shallow architecture with a single hidden layer. The input and hidden layer widths are $[d_{\mathrm{x}},128]$, where $d_{\mathrm{x}}$ is the dimensionality of the input. The output layer is composed of three heads, corresponding to the mean, standard deviation and covariance outputs. The mean and standard deviation heads have output sizes of $d_{\mathrm{z}}$, while the covariance head has an output size of $d_{\mathrm{z}}^{2}$, where $d_{\mathrm{z}}=d_{\mathrm{z_{\mathrm{x}}}}+d_{\mathrm{z_{\mathrm{c}}}}+d_{\mathrm{z_{\mathrm{y}}}}$, with $d_{\mathrm{z_{\mathrm{i}}}}$ denoting the size of the $i$’th subset of the latent space. A ReLU activation function is applied on all layers except the final output layer. For the introductory examples, an independent encoder network is used for each subset of the latent variables. The hidden layer widths of each independent network are set to 64 units, and the output shapes are adjusted according to the dimensionality of the corresponding subset of the latent variables. To further ensure numerical stability, the outputs of all encoder NNs are clamped within ranges of values that are expected to be encountered for the case studies investigated in this work.
@@
Internal workspace
To facilitate the application of the model to cases involving physics-grounded latent variables with bounded support, and also to improve numerical stability, all parameters are transformed from an unbounded and normalized internal workspace to the original space in which they are defined. This is achieved by applying a sequence of deterministic transformations to the samples and corresponding scaling of the log-densities. In the following, variables in the internal workspace are denoted as $u$. The samples at the output of the encoder are first bounded by applying the logistic transform $u^{\prime}=\frac{1}{1+e^{-u}}$, and subsequently scaled and shifted using an affine transform $z=u^{\prime}\cdot(\mathrm{UB}_{\mathrm{z}}-\mathrm{LB_{\mathrm{z}}})+\mathrm{LB}_{\mathrm{z}}$ to bound the variables to their specified supports defined by the lower and upper bound $\mathrm{LB}$ and $\mathrm{UB}$. The samples and densities can also be mapped from the original space to the internal workspace by applying the corresponding inverse transforms in reverse order.
@@
Optimization is carried out using the Adam algorithm [Kingma2017] with minibatch gradient estimation [Kingma2022]. The model is trained for up to $20,000$ iterations with a batch size of $64$. Early stopping is implemented by monitoring the value of the ELBO, evaluated on a held-out validation set with size $N_{\mathrm{val}}=512$. The training is terminated if no improvement of the ELBO is observed over $2,000$ iterations. Gradient and objective estimates are obtained using $16$ Monte Carlo samples during training, $64$ during validation, and $512$ during evaluation, although in practice the training was found to be insensitive to the number of samples. All learning rates are set to $0.001$, except for the learning rate of the standard deviation parameter for the response $\sigma_{\mathrm{x}}$, which is set to $0.005$.
Plotting

### v3

@@ 6.2 Assumptions and limitations
It is important to consider the assumptions and limitations of the proposed approach. One of the main drawbacks of the model is the additivity assumption imposed on the physical, domain and class components of the response. It is expected that the model will perform sub-optimally when this assumption is violated. Furthermore, the accuracy of the inferred physics-grounded latent variables will depend on the relative contribution of the physics, domain and class influences to the measured response. Significant domain and class contributions to the response, or violating the additivity assumption of Equation 2.3, can lead to inaccurate inference of physics-grounded latent variables and large uncertainty in the predictions. Additionally, the model requires that multiple types of data are available, namely measurements of the structural response and information on domain and class. In SHM applications, this might necessitate data alignment procedures of response measurements, environmental conditions and damage level descriptions, and could potentially limit the immediate applicability of the proposed architecture. It is worth mentioning that the interaction between the encoder, decoders, the GRL, and the known physics can be unintuitive in some applications, limiting the applicability of the approach and potentially necessitating implicit supervision by a human expert.
6.3 Practical considerations
Specifying an appropriate value of $\lambda$ for a given learning problem is not straightforward. Schemes for scheduling or adaptively tuning the strength of the GRL during training have been proposed [55, 45, 25], but have not been considered in this work. Instead, we focus on providing intuition and clarity regarding the influence of $\lambda$ through the qualitative and quantitative results presented in Section 5. Furthermore, it is known that adversarial training can be unstable [73]. Throughout this work, occasional instability and overfitting were observed when using small datasets and large batch sizes. Depending on the case study and the value of the $\lambda$ hyperparameter, oscillatory behavior may also occur. We found that these issues could be addressed by adjusting the $\lambda$ hyperparameter, implementing early stopping based on the value of the ELBO on a held-out validation set, and reducing the batch size.
The dimensionality of the latent space is an important design parameter in VAE, and excessively small or large dimensionality can result in poor reconstruction quality [21]. Depending on the available computational budget and problem complexity, approaches for determining an appropriate dimensionality are often based on manual trial and error or grid search [60]. More sophisticated approaches include dynamically adjusting the number of latent variables during optimization [18, 60], automatic relevance determination [59] and multi-stage models [17]. A key advantage of VAE in engineering, physical, and scientific applications, is that domain knowledge can guide reasoning about the type and number of the dominant generative factors in the data, informing the design of the latent space. VAE are generally insensitive to over-specification of the latent space dimensionality, with superfluous dimensions becoming inactive and ignored by the decoder [74, 4]. Choosing the dimensionality of the domain and class latent space to be a multiple of the expected number of generative factors, based on domain knowledge, and subsequently refining this choice by monitoring the number of inactive dimensions after training can therefore be a viable approach.
It is not possible to provide a rule-of-thumb about the amount of data required for effective training. This would be dependent on the specific problem, noise levels, physics-based model, and accuracy of the domain and class information, and also on the particular architecture choices (e.g. the number and depth of layers used in the feed-forward NNs in the encoder and decoder). In some applications, the incorporation of the known physics might lead to a reduction in the data requirements. This has not been investigated in the current paper since it is believed that it would be strongly dependent on the case study chosen, rather than offering any general insight.
@@ 7 Conclusions
The present work contributes to the emerging applications of probabilistic generative models in engineering, by investigating disentangled and invariant representation learning as a tool for grounding VAE to the known physics. Specifically, a Physics-Enhanced Machine Learning strategy utilizing a VAE architecture is proposed, with the aim of learning a disentangled representation of physical, domain and class confounding influences that are present in the response measurements of physical systems. This is achieved by having the decoder and latent space of the VAE be semantically and functionally separated into data-driven and physics-grounded branches. An easy to implement regularization method based on the GRL is used to constrain the data-driven components, resulting in a model that preferentially utilizes the known physics. An interpretable and intuitive hyperparameter is used to specify the strength of GRL, and wether the model is trained in a collaborative or adversarial manner. Moreover, a strategy for quantifying the type and relative amount of information encoded in different sets of latent variables is proposed, yielding insights on the degree of disentanglement achieved by the model.
@@
The code and data (generated via the synthetic use cases) needed to replicate the results shown in this paper can be accessed through the link: https://doi.org/10.5281/zenodo.15813028 [43].
@@ Appendix A Implementation details
The encoder reduces the dimensionality of the input measurements and maps them to vectors of mean values $\bm{\mu}_{\mathrm{\phi}}(\bm{x})$, standard deviations $\bm{\sigma}_{\mathrm{\phi}}(\bm{x})$ and a lower triangular matrix $\bm{L}^{\prime}_{\mathrm{\phi}}(\bm{x})$. The lower triangular factor $\bm{L}_{\mathrm{\phi}}(\bm{x})=\bm{L}^{\prime}_{\mathrm{\phi}}(x)+\bm{\sigma}_{\mathrm{\phi}}(\bm{x})\bm{I}$ is the Cholesky decomposition factor of the covariance matrix $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})$, i.e. $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})=\bm{L_{\mathrm{\phi}}(\bm{x})L_{\mathrm{\phi}}(\bm{x})}^{T}$, such that the posterior distribution corresponding to each input is a multivariate Normal distribution. The encoder outputs the log of the standard deviations, which are then exponentiated to avoid negative values. The reparametrization trick [41] is exploited to define the latent variables $\bm{z}$ as a deterministic transformation of a noise variable $\bm{\epsilon}\sim p(\bm{\epsilon})$. This facilitates the computation of unbiased Monte Carlo gradient estimates of the objective with respect to the variational parameters, using automatic differentiation. With the exception of the introductory examples presented in Section 2.4, the encoder is everywhere formulated as a single feed-forward NN using a shallow architecture with a single hidden layer. The input and hidden layer widths are $d_{\mathrm{x}},128]$, where $d_{\mathrm{x}}$ is the dimensionality of the input. The output layer is composed of three heads, corresponding to the mean, standard deviation and covariance outputs. The mean and standard deviation heads have output sizes of $d_{\mathrm{z}}$, while the covariance head has an output size of $d_{\mathrm{z}}^{2}$, where $d_{\mathrm{z}}=d_{\mathrm{z_{\mathrm{x}}}}+d_{\mathrm{z_{\mathrm{c}}}}+d_{\mathrm{z_{\mathrm{y}}}}$, with $d_{\mathrm{z_{\mathrm{i}}}}$ denoting the size of the $i$’th subset of the latent space. A ReLU activation function is applied on all layers except the final output layer. For the introductory examples, an independent encoder network is used for each subset of the latent variables. The hidden layer widths of each independent network are set to 64 units, and the output shapes are adjusted according to the dimensionality of the corresponding subset of the latent variables. To further ensure numerical stability, the outputs of all encoder NNs are clamped within ranges of values that are expected to be encountered for the case studies investigated in this work.
@@
Latent variable transformation
To facilitate the application of the model to cases involving physics-grounded latent variables with bounded support, and to improve numerical stability, all parameters are transformed from an unbounded and normalized base latent space to the target latent space in which they are defined. This is achieved by applying a sequence of deterministic transformations to the samples and corresponding scaling of the log-densities. In the following, variables in the base space are denoted as $u$. The samples at the output of the encoder are first bounded by applying the logistic transform $u^{\prime}=\frac{1}{1+e^{-u}}$, and subsequently scaled and shifted using an affine transform $z=u^{\prime}\cdot(\mathrm{UB}-\mathrm{LB})+\mathrm{LB}$ to bound the variables to their specified supports defined by the lower and upper bound $\mathrm{LB}$ and $\mathrm{UB}$. The samples and densities can also be mapped from the target latent space to the base latent space by applying the corresponding inverse transforms in reverse order.
@@
Optimization is carried out using the Adam algorithm [39] with minibatch gradient estimation [41]. The model is trained for up to $20,000$ iterations with a batch size of $64$. Early stopping is implemented by monitoring the value of the ELBO, evaluated on a held-out validation set with size $N_{\mathrm{val}}=512$. The training is terminated if no improvement of the ELBO is observed over $2,000$ iterations. Gradient and objective estimates are obtained using $16$ Monte Carlo samples during training, $64$ during validation, and $512$ during evaluation, although in practice the training was found to be insensitive to the number of samples. All learning rates are set to $0.001$, except for the learning rate of the standard deviation parameter for the response $\sigma_{\mathrm{x}}$, which is set to $0.005$. The $\alpha$ and $\beta$ hyperparameters of the optimization objective are taken as $\beta=\alpha_{\mathrm{x}}=\alpha_{\mathrm{c}}=\alpha_{\mathrm{y}}=1.0$ for all the experiments presented in this work.
Visualization